marginal cost is 20 and the price elasticity of demand is -2.0 determine profit maximizing price


1
Expert's answer
2021-02-04T07:52:52-0500

"Price=MC\times[e/(1+e)]"

"where,MC=\$20, e=-2.0"

so price "=\$20\times[(-2.0)/(1+(-2.0))]=20\times[(-2.0)/(-1)]=20\times2=\$40"

So the profit maximizing price "=\$40"